OpenStudy (anonymous):

Please, this is the 2nd to last question I need! I will fail if you don't answer!

OpenStudy (anonymous):

10.07] Trevor is tiling his bathroom floor, which has an area that is represented as 175r5 square inches. Each tile has an area of one hundred seventy five r to the fifth power, all over the square root of the quantity sixty four r to the seventeenth power. The total number of tiles used can be represented by the expression below. one hundred seventy five r to the fifth power, all over the square root of the quantity sixty four r to the seventeenth power Simplify the expression for the total number of tiles used. Show your work.

OpenStudy (anonymous):

is it \[\frac{125r^5}{\sqrt{64r^7}}\]?

OpenStudy (anonymous):

For which one the 1st problem?

OpenStudy (anonymous):

\[\frac{125r^5}{\sqrt{64r^{17}}}\]

OpenStudy (anonymous):

\[\sqrt{64r^{17}}=8r^8\sqrt{r}\]

OpenStudy (anonymous):